OpenStudy (anonymous):

A 800MW electric power plant has an efficiency of 30%. The energy which is not useful output becomes waste heat. How much waste heat is discharged by this plant every second?

OpenStudy (anonymous):

30% = 800Mw/Pin, correct? But I'm not sure how to get from W to J...

OpenStudy (anonymous):

1 watt is 1 joule per second

OpenStudy (anonymous):

so 800MW = 800Mj per second

OpenStudy (anonymous):

But when I solved for 30% = 800MW/Pin, I got 2666 MW, which is decidedly not correct (the right answer is 1800 MJ)

OpenStudy (anonymous):

70% is the waste energy, 70% of 800MW is 560MW = 560Mj

OpenStudy (anonymous):

how can the plant produce more waste heat than heat going in? 1800Mj > than 800Mj what is "Pin"?

OpenStudy (anonymous):

The amount of power being produced. 800MW is the output of power.

OpenStudy (anonymous):

0.3*power in = 800 power in = 800/0.3 = 2667 excess heat = power in * 0.7 = 2667*0.7 = 1866

OpenStudy (anonymous):

Ah, okay. Thank you.